Angela needs 20 quarts of 50% antifreeze solution in her radiator. She plans to obtain this by mixing some pure antifreeze with an appropriate amount of a 40% antifreeze solution. How many quarts of each should she use?
Answer by ankor@dixie-net.com(22740)

Angela needs 20 quarts of 50% antifreeze solution in her radiator. She plans to obtain this by mixing some pure antifreeze with an appropriate amount of a 40% antifreeze solution. How many quarts of each should she use?
:
Let x = amt of pure anti-freeze required
Then
(20-x) = amt of 40% solution required:
:
The percent equation:
.40(20-x) + 1.0(x) = .50(20)
8 - .4x + x = 10
.6x = 10 - 8
.6x = 2
x = 2/.6
x = 3.33 qt of pure anti freeze
and
20 - 3.33 = 16.67 qt of 40% solution
:
:
Check our solution
.4(16.67) + 3.33 = .5(20)
6.668 + 3.33 = 9.998 ~ 10
: